# ObjectVLA: End-to-End Open-World Object Manipulation Without Demonstration

### v1

@@ 5 Conclusion
In this work, we present ObjectVLA, a Vision-Language-Action framework that addresses object generalization in robotic manipulation. By integrating vision-language datasets with robot interaction data, our method establishes a unified pipeline that bridges semantic understanding and physical action execution. This enables zero-shot generalization to over 100 novel objects with a 64% success rate, even when objects differ in category, appearance, or fine-grained attributes (e.g., color, shape). Our framework demonstrates that lightweight co-training with visual-textual priors and localization-aware reasoning can unlock robust cross-modal alignment. Key to our success is the ability to adapt rapidly to real-world scenarios: using just a few smartphone-captured images and quick continual finetuning, robots generalize to unseen objects without costly human demonstrations. We validate our approach across diverse tasks—including bin-picking, rotating and pushing—showcasing its versatility and practicality. Our results highlight a path toward scalable robotic learning systems that reduce dependence on large-scale teleoperation data while maintaining high performance.
@@ 6 Limitation
There are still a number of limitations in this work. Specifically, the visual-text data was collected by the authors using either a robot-mounted camera or a smartphone. While we have not yet explored the feasibility of leveraging internet-sourced visual-text data, it presents an intriguing avenue for future research. Specifically, investigating the necessary degree of visual similarity between internet images and target objects for effective skill transfer would be valuable. Our primary focus here is to introduce a novel pipeline that enables deep learning models to transfer skills to new objects without explicit demonstrations. Determining the limits of this transferability, particularly concerning the permissible visual gap between training and target objects, remains an open question for future investigation. Currently, our method struggles to generalize to novel backgrounds and lighting conditions. We believe the visual gap between our collected visual-text data and the robot’s operational environment contributes to this challenge. Bridging this gap to improve generalization is a key focus for future development.

### v2

@@ 5 Conclusion
In this work, we present ObjectVLA, a Vision-Language-Action framework that addresses object generalization in robotic manipulation. By integrating vision-language datasets with robot interaction data, our method establishes a unified pipeline that bridges semantic understanding and physical action execution. This enables zero-shot generalization to over 100 novel objects with a 64% success rate, even when objects differ in category, appearance, or fine-grained attributes (e.g., color, shape). Our framework demonstrates that lightweight co-training with image-text priors and localization-aware reasoning can unlock robust cross-modal alignment. Key to our success is the ability to adapt rapidly to real-world scenarios: using just a few smartphone-captured images and quick continual finetuning, robots generalize to unseen objects without costly human demonstrations. We validate our approach across diverse tasks—including bin-picking, rotating and pushing—showcasing its versatility and practicality. Our results highlight a path toward scalable robotic learning systems that reduce dependence on large-scale teleoperation data while maintaining high performance.
@@ 6 Limitation
There are still a number of limitations in this work. Specifically, the image-text data was collected by the authors using either a robot-mounted camera or a smartphone. While we have not yet explored the feasibility of leveraging internet-sourced image-text data, it presents an intriguing avenue for future research. Specifically, investigating the necessary degree of visual similarity between internet images and target objects for effective skill transfer would be valuable. Our primary focus here is to introduce a novel pipeline that enables deep learning models to transfer skills to new objects without explicit demonstrations. Determining the limits of this transferability, particularly concerning the permissible visual gap between training and target objects, remains an open question for future investigation. Currently, our method struggles to generalize to novel backgrounds and lighting conditions. We believe the visual gap between our collected image-text data and the robot’s operational environment contributes to this challenge. Bridging this gap to improve generalization is a key focus for future development.